A data set lists earthquake depths. The summary statistics are n =500 , x =6.78km, s= 4.41 km. Use a 0.01 significance level to test the claim of a seismologist that these earthquakes are from a population with a mean equal to 6.00. Assume that a simple random sample has been selected. Identify the null and alternative hypotheses, test statistic, P-value, and state the final conclusion that addresses the original claim.

What are the null and alternative hypotheses?

A. H0 : μ ≠ 6.00km H1 : μ = 6.00km

B. H0 : μ =6.00km H1 : μ >6.00km

C. H0 : μ =6.00km H1 : μ <6.00km

D. H0 : μ =6.00km H1 : μ ≠ 6.00km

Determine the test statistic.

(Round to two decimal places as needed.)

Determine the P-value.

(Round to three decimal places as needed.)

State the final conclusion that addresses the original claim.

▼ Reject Fail to reject

H0.

There is

▼ not sufficient sufficient

evidence to conclude that the original claim that the mean of the population of earthquake depths is 6.00km

▼is not is correct.
Answer #1

To Test :-

D. H0 : μ =6.00km

H1 : μ ≠ 6.00km

Test Statistic :-
t = ( X̅ - µ ) / ( S / √(n))
t = ( 6.78 - 6 ) / ( 4.41 / √(500) )
t = 3.96

Test Criteria :-
Reject null hypothesis if | t | > t(α/2, n-1)
t(α/2, n-1) = t(0.01 /2, 500-1) = 2.586
| t | > t(α/2, n-1) = 3.955 > 2.586
Result :- Reject null hypothesis

Decision based on P value
P - value = P ( t > 3.955 ) = 0.0001
Reject null hypothesis if P value < α = 0.01 level of significance
P - value = 0.0001 < 0.01 ,hence we reject null hypothesis
Conclusion :- Reject null hypothesis

There is not sufficient evidence to conclude that the original claim that the mean of the population of earthquake depths is 6.00 km.